Consider a beaker containing equal amounts of dihydrogen phosphate anhd hydrogen phosphate to which either a strong acid is added (Beaker C) or a strong base is added (beaker D).

1) Calculate the pH of the contents of beaker C

2) Calculate the pH of the contents of beaker D.

256 Activity 51 l What Are Buffer systems in Your Body that Keep You Alive phos Consider a beaker the dilhydrogen phosphate and hydrogen containing equal amounts of strong base is added phate to which either a strong acid i added 1.0 mL 1.0 mL 1.0 M NaOH 1.0 M HCI 1.0 L 1.0 L [H2PO4 LH2PO41 0.010 M 0.010 M UHPO42 0.010 M 0.010 M Beaker D Beaker C 9. Calculate the pH of the contents of Beaker C
Answer #1

beaker C

moles of H2PO4- = 0.01

moles of HPO42- = 0.01

moles of acid added = 1 x 1 / 1000 = 0.001

pH = pKa + log [salt - C / acid + C]

    = 7.21 + log [0.01 - 0.001 / 0.01 + 0.001]

    = 7.12

pH = 7.12

beaker D

moles of base added = 1 x 1 / 1000 = 0.001

pH = pKa + log [salt + C / acid - C]

    = 7.21 + log [0.01 + 0.001 / 0.01 - 0.001]

   = 7.30

pH = 7.30